(1 point) Suppose h(x) = f(x) and the equation of the tangent line to f(x) at x = 1 Is y = 4+2(x - 1). Find h (1) h(1) =
Answer #1

Solution => Given that 0 h(a) = dfca) tangent and the es 2=1 line to equation of the fins at f(x)= y = = 4+2(2-1) f(x) = 4+2(